# Beyond One-Hot Labels: Semantic Mixing for Model Calibration

### v1

@@ Abstract
Model calibration seeks to ensure that models produce confidence scores that accurately reflect the true likelihood of their predictions being correct. However, existing calibration approaches are fundamentally tied to datasets of one-hot labels implicitly assuming full certainty in all the annotations. Such datasets are effective for classification but provides insufficient knowledge of uncertainty for model calibration, necessitating the curation of datasets with numerically rich ground-truth confidence values. However, due to the scarcity of uncertain visual examples, such samples are not easily available as real datasets. In this paper, we introduce calibration-aware data augmentation to create synthetic datasets of diverse samples and their ground-truth uncertainty. Specifically, we present Calibration-aware Semantic Mixing (CSM), a novel framework that generates training samples with mixed class characteristics and annotates them with distinct confidence scores via diffusion models. Based on this framework, we propose calibrated reannotation to tackle the misalignment between the annotated confidence score and the mixing ratio during the diffusion reverse process. Besides, we explore the loss functions that better fit the new data representation paradigm. Experimental results demonstrate that CSM achieves superior calibration compared to the state-of-the-art calibration approaches. Code is available at github.com/E-Galois/CSM.
@@ 1 Introduction
To generate samples with ground-truth uncertainty, a naive solution could be Mixup-based approaches (Yun et al., 2019; Kim et al., 2020; Uddin et al., 2020; Chen et al., 2022; Noh et al., 2023). Through augmenting training data with pixel-interpolated samples and soft labels, DNNs are supposed to learn diverse confidence. However, such methods often produce low-fidelity samples due to unrealistic pixel fusion or fragmentary collage, which deviate significantly from the distribution of real-world data as shown in Figure 1. Consequently, these approaches struggle to generalize the learned calibration knowledge to real-world scenarios and hardly contribute to the calibration improvement (Wen et al., 2020; Maroñas et al., 2021; Wang et al., 2023a).
@@
Besides these image samples, their corresponding soft labels should be annotated for learning. A naive solution is the mixing ratio during the diffusion reverse process. However, it can rarely indicate the class posterior precisely. To tackle this issue, we propose to reannotate the data by leveraging CLIP’s visual features which have been jointly trained with massive language supervision (Radford et al., 2021), and are thus able to provide more precise class posteriors. By representing the optimized sample embeddings as interpolations of class prototypes, we can identify and eliminate class-specific biases in the estimated mixing coefficients, resulting in more accurate and effective annotations. Moreover, we highlight the problem of imbalanced fitting of augmented data when traditional loss functions for calibration are utilized through some theoretical analysis. We reveal that the training of augmented data with $\mathcal{L}_{2}$ loss intrinsically results in balanced learning and better calibration performance. Extensive experiments on various benchmarks and tasks demonstrate our method’s strong effectiveness in accurately estimating confidence levels, achieving superior results for data-aware model calibration.
@@ 2.1 Model Calibration
Network calibration (Guo et al., 2017) requires techniques to align model confidence with actual accuracy. Post-hoc methods adjust test-time parameters using a validation set to improve calibration. Notable techniques include Temperature Scaling (TS) (Guo et al., 2017), Histogram Binning (Zadrozny & Elkan, 2001), Beta calibration (Kull et al., 2017) and its extension to Dirichlet calibration (Kull et al., 2019). While simple, post-hoc methods are sensitive to distributional shifts (Ovadia et al., 2019). Training-time regularization methods involve explicit learning constraints and implicit loss functions. Label Smoothing (LS) (Müller et al., 2019) reduces output entropy by replacing one-hot targets with smoothened labels. Margin-based Label Smoothing (MbLS) (Liu et al., 2022) balances model discriminability and calibration by imposing a margin on logits. Implicit loss functions, such as those optimizing Expected Calibration Error (Karandikar et al., 2021; Kumar et al., 2018), Focal loss (Ross & Dollár, 2017; Mukhoti et al., 2020), Inverse Focal loss (Wang et al., 2021), and Mean Square Error (Hui & Belkin, 2020; Liang et al., 2024), are effective objectives to improve calibration. Data-aware calibration primarily involves perturbation methods (Tao et al., 2024) and Mixup-based methods (Zhang et al., 2018; Yun et al., 2019; Hendrycks et al., 2019), originally designed for improved generalization . These techniques have been found effective for calibration (Zhang et al., 2018; Pinto et al., 2022). RegMixup (Pinto et al., 2022) uses Mixup as a regularizer for cross-entropy loss, enhancing both accuracy and uncertainty estimation. RankMixup (Noh et al., 2023) incorporates ordinal ranking relationships among samples to reduce Mixup label bias. However, recent work (Wang et al., 2023a) highlights the limitation of simple convex combinations in Mixup methods.
@@ 3.2 Calibration-aware Data Augmentation
A straightforward calibration scheme generating soft-labeled data is the mixup (Zhang et al., 2018) augmentation, which generates fused samples and labels from random training pairs as $\widetilde{\bm{x}}=\lambda\bm{x}_{i}+(1-\lambda)\bm{x}_{j}$ and $\widetilde{\bm{q}}=\lambda\bm{q}_{i}+(1-\lambda)\bm{q}_{j}$, where $\lambda\sim\operatorname{Beta}(\alpha,\alpha)\in[0,1]$ is a mixup coefficient, $\alpha>0$ is a hyperparameter controlling the interpolation strength (Noh et al., 2023), and $\bm{q}_{i}$ is the one-hot vector of $y_{i}$. However, the direct utilization of mixup cannot contribute to the calibration improvement, as empirically found by several studies (Wen et al., 2020; Maroñas et al., 2021; Wang et al., 2023a). We mainly attribute this failure to the fact that mixup performs calibration learning on low-fidelity images due to simple convex combinations, which makes the learned calibration difficult to generalize to real-world data. Therefore, to establish an effective calibration-aware data augmentation, we propose to generate high-fidelity images with semantic mixing, which reflects the shifting process of class posteriors from one category to another.
We start by expressing the general problem as how to decide and sample from vicinal distribution
@@
To blend class semantics with soft labels, we propose a dual-conditioned sampling strategy via a conditional diffusion model, i.e., $p_{\phi}(\bm{x}\mid\bm{q}^{y})=\int_{\bm{x}^{(T)}}p_{\phi}(\bm{x}\mid\bm{q}^{y},\bm{x}^{(T)})p(\bm{x}^{(T)})\mathrm{d}\bm{x}^{(T)}$ which is explicitly optimized to estimate $p(\bm{x}\mid y)$. Specifically, we generate augmented samples $\{\widetilde{\bm{x}}^{\hat{\lambda}}\}$ via diffusion reverse process as
@@ 3.3 Calibrated Reannotation and Balanced Learning
See Section A.1 for proof. It is evident that directly adopting Eq. (9) to calculate $\lambda$ would induce bias due to the unequal class norms in $\bm{P}_{i}^{2}-\bm{P}_{j}^{2}\neq 0$ and inter-class distances in $(\bm{P}_{i}-\bm{P}_{j})^{2}\neq(\bm{P}_{k}-\bm{P}_{l})^{2}$ for $i\neq j$ and $\{i,j\}\neq\{k,l\}$. Note that $\lambda_{E}$ is only measured by the affine space of the class prototypes regardless of these biases. We propose to regard these factors as invariant across classes and simplify Eq. (9) as
@@
The formal proofs of Propositions 3.2, 3.3, and 3.4 are provided in Appendix A.2. Notably, this objective is tailored for learning on soft-labeled data, in contrast to the calibration of one-hot-labeled samples explored in (Liang et al., 2024). In this way, we can formulate our overall objective as a simple combination of the CE and L2 losses, i.e.,
@@ 4.1 Results
We also provide qualitative comparisons using Wide-ResNet-26-10 (Zagoruyko, 2016) and DenseNet-121 (Huang et al., 2017) as the network in Appendix C, which demonstrate similar properties to those in Table 1. Meanwhile, we provide results on the Eq-Data setting in Table 2 for comparisons with non-augmentation methods for equalized training time. It is observed that our CSM can still achieve competitive or superior performance, particularly in consistently superior post-temperature ECE and AECE.
@@ 4.2 Discussions
w/o Augmented Data 77.48 12.94 12.94 2.63 2.66 1.6
Mixup (CE) 78.47 8.68 8.68 2.14 2.19 1.3
Mixup (FL) 79.64 2.45 2.44 2.45 2.44 1.0
Mixup (L2) 78.88 2.49 2.62 2.49 2.62 1.0
CSM (CE) 78.93 2.47 2.14 2.47 2.14 1.0
CSM (FL) 77.65 2.43 2.30 2.16 2.09 0.9
CSM (L2) w/o reanno. 78.92 5.20 5.20 3.33 3.24 1.2
CSM (L2) 78.84 1.29 1.63 1.29 1.63 1.0
We compare our CSM with variants in Table 4, testing different augmentation types, soft labels, and loss functions. We can observe that Mixup and CSM variants outperform the standard CE loss, improving both ECE and accuracy, highlighting the importance of data-driven network calibration. Among the losses used in CSM, the L2 loss is the only one that reduces ECE, AECE, and post-TS results below 2.0, suggesting it effectively balances learned confidence in augmentations. In contrast, CE and FL losses often require temperature adjustments, with CE favoring sharper labels and FL for softer ones, aligning with our theoretical expectations from Section 3.3. CSM variants, particularly with CE and L2 losses, significantly outperform traditional Mixup in terms of calibration, verifying the effectiveness of CSM’s realistic in-domain augmentations.
However, applying L2 loss to the traditional Mixup samples does not yield significantly better results. It is noteworthy that CSM uses conditioned dual-sampling of one-hot samples $\widetilde{\bm{x}}^{0}$ and $\widetilde{\bm{x}}^{1}$, which are correlated considering joint latent variable $\bm{x}^{(T)}$, while every pair of them in Mixup is sampled independently, which makes the soft-labeled samples much easier to be distinguished apart by the model. Therefore, traditional Mixup with L2 loss cannot perform noticeably better on all metrics.
Finally, compared with the variant w/o reannotation (or using the labels in generation), our reannotated variant produces significantly better calibration performances, validating the accurate confidence estimation in our refined labels.
@@
Augmented Samples & Annotations To analyze the properties of augmented samples, we visualize some case images in Fig. 3. It can be seen that CSM can produce semantically sound intermediate samples compared to Mixup’s simple overlap. Meanwhile, these samples exhibit characteristics of both classes with gradual semantic change from one category to another. For images with abrupt transitions (in red numbers) or inaccurate class posteriors, our annotation paradigm yields more precise $\lambda$ values compared to the generation phase $\hat{\lambda}$.
Reliability Diagrams and Confidence Values To analyze the learned confidence values, we compare the reliability diagrams of different methods across Figure 4. We examine CSM alongside three variants: 1) without augmentations, 2) with vanilla Mixup, and 3) using labels from generation. The augmentation-free variant shows severe over-confidence across the middle-to-high confidence range, while the Mixup variant also exhibits over-confidence, though less severely. Regularization-based methods that calibrate with one-hot labels can show similar over-confidence, highlighting common issues in confidence estimation. CSM with primitive labels $\hat{\lambda}$, however, is both over- and under-confident at higher and lower ranges, respectively, suggesting a misalignment between annotations and confidence levels. In contrast, CSM demonstrates precise confidence estimation across both low and high ranges, with only slight errors in the middle, confirming the effectiveness of our confidence-aware augmentation.

### v2

@@ Abstract
Model calibration seeks to ensure that models produce confidence scores that accurately reflect the true likelihood of their predictions being correct. However, existing calibration approaches are fundamentally tied to datasets of one-hot labels implicitly assuming full certainty in all the annotations. Such datasets are effective for classification but provides insufficient knowledge of uncertainty for model calibration, necessitating the curation of datasets with numerically rich ground-truth confidence values. However, due to the scarcity of uncertain visual examples, such samples are not easily available as real datasets. In this paper, we introduce calibration-aware data augmentation to create synthetic datasets of diverse samples and their ground-truth uncertainty. Specifically, we present Calibration-aware Semantic Mixing (CSM), a novel framework that generates training samples with mixed class characteristics and annotates them with distinct confidence scores via diffusion models. Based on this framework, we propose calibrated reannotation to tackle the misalignment between the annotated confidence score and the mixing ratio during the diffusion reverse process. Besides, we explore the loss functions that better fit the new data representation paradigm. Experimental results demonstrate that CSM achieves superior calibration compared to the state-of-the-art calibration approaches. Our code is available here.
@@ 1 Introduction
To generate samples with ground-truth uncertainty, a naive solution could be Mixup-based approaches (Yun et al., 2019; Kim et al., 2020; Uddin et al., 2021; Chen et al., 2022; Noh et al., 2023). By augmenting training data with pixel-interpolated samples and soft labels, DNNs are supposed to learn diverse confidence. However, such methods often produce low-fidelity samples due to unrealistic pixel fusion or fragmentary collage, which deviate significantly from the distribution of real-world data as shown in Figure 1. Consequently, these approaches struggle to generalize the learned calibration knowledge to real-world scenarios and hardly contribute to the calibration improvement (Wen et al., 2021; Maroñas et al., 2021; Wang et al., 2023a). Data augmentations have also been explored at test time (Hekler et al., 2023) to improve uncertainty estimation. While these post-hoc methods can mitigate over-confidence globally, they often fail to achieve precise calibration because they ignore miscalibration in the training with one-hot samples.
@@
Besides these image samples, their corresponding soft labels should be annotated for learning. A naive solution is the mixing ratio during the diffusion reverse process. However, it can rarely indicate the class posterior precisely. To tackle this issue, we propose to reannotate the data by leveraging CLIP’s visual features which have been jointly trained with massive language supervision (Radford et al., 2021), and are thus able to provide more precise class posteriors. By representing the optimized sample embeddings as interpolations of class prototypes, we can identify and eliminate class-specific biases in the estimated mixing coefficients, resulting in more accurate and effective annotations. Moreover, we highlight the problem of imbalanced fitting of augmented data when utilizing traditional loss functions for calibration through some theoretical analysis. We reveal that the training of augmented data with $\mathcal{L}_{2}$ loss intrinsically results in balanced learning and better calibration performance. Extensive experiments on various benchmarks and tasks demonstrate our method’s strong effectiveness in accurately estimating confidence levels, achieving superior results for data-aware model calibration.
@@ 2.1 Model Calibration
Network calibration (Guo et al., 2017) requires techniques to align model confidence with actual accuracy. Post-hoc methods adjust test-time parameters using a validation set to improve calibration. Notable techniques include Temperature Scaling (TS) (Guo et al., 2017), Histogram Binning (Zadrozny & Elkan, 2001), Beta calibration (Kull et al., 2017) and its extension to Dirichlet calibration (Kull et al., 2019). While simple, post-hoc methods are sensitive to distributional shifts (Ovadia et al., 2019). Training-time regularization methods involve explicit learning constraints and implicit loss functions. Label Smoothing (LS) (Müller et al., 2019) reduces output entropy by replacing one-hot targets with smoothened labels. Margin-based Label Smoothing (MbLS) (Liu et al., 2022) balances model discriminability and calibration by imposing a margin on logits. Implicit loss functions, such as those optimizing Expected Calibration Error (Karandikar et al., 2021; Kumar et al., 2018), Focal loss (Ross & Dollár, 2017; Mukhoti et al., 2020), Inverse Focal loss (Wang et al., 2021), and Mean Square Error (Hui & Belkin, 2021; Liang et al., 2024), are effective objectives to improve calibration. Data-aware calibration primarily involves perturbation methods (Tao et al., 2024) and Mixup-based methods (Zhang et al., 2018; Yun et al., 2019; Hendrycks et al., 2019), originally designed for improved generalization . These techniques have been found effective for calibration (Zhang et al., 2018; Pinto et al., 2022). RegMixup (Pinto et al., 2022) uses Mixup as a regularizer for cross-entropy loss, enhancing both accuracy and uncertainty estimation. RankMixup (Noh et al., 2023) incorporates ordinal ranking relationships among samples to reduce Mixup label bias. However, recent work (Wang et al., 2023a) highlights the limitation of simple convex combinations in Mixup methods.
@@ 3.2 Calibration-aware Data Augmentation
A straightforward calibration scheme generating soft-labeled data is the mixup (Zhang et al., 2018) augmentation, which generates fused samples and labels from random training pairs as $\widetilde{\bm{x}}=\lambda\bm{x}_{i}+(1-\lambda)\bm{x}_{j}$ and $\widetilde{\bm{q}}=\lambda\bm{q}_{i}+(1-\lambda)\bm{q}_{j}$, where $\lambda\sim\operatorname{Beta}(\alpha,\alpha)\in[0,1]$ is a mixup coefficient, $\alpha>0$ is a hyperparameter controlling the interpolation strength (Noh et al., 2023), and $\bm{q}_{i}$ is the one-hot vector of $y_{i}$. However, the direct utilization of mixup cannot contribute to the calibration improvement, as empirically found by several studies (Wen et al., 2021; Maroñas et al., 2021; Wang et al., 2023a). We mainly attribute this failure to the fact that mixup performs calibration learning on low-fidelity images due to simple convex combinations, which makes the learned calibration difficult to generalize to real-world data. Therefore, to establish an effective calibration-aware data augmentation, we propose to generate high-fidelity images with semantic mixing, which reflects the shifting process of class posteriors from one category to another.
We start by expressing the general problem as how to establish a vicinal distribution that can be sampled from, i.e.,
@@
To blend class semantics with soft labels, we propose a dual-conditioned sampling strategy via a conditional diffusion model explicitly optimized to estimate $p(\bm{x}\mid y)$, i.e.,
$p_{\phi}(\bm{x}\mid\bm{q}^{y})=\int_{\bm{x}^{(T)}}p_{\phi}(\bm{x}\mid\bm{q}^{y},\bm{x}^{(T)})p(\bm{x}^{(T)})\mathrm{d}\bm{x}^{(T)},$ (5)
Specifically, we generate augmented samples $\{\widetilde{\bm{x}}^{\hat{\lambda}}\}$ via diffusion reverse process as
@@ 3.3 Calibrated Reannotation and Balanced Learning
See Section A.1 for proof of the above Equations. It is evident that directly adopting Eq. (10) to calculate $\lambda$ would induce bias due to the unequal class norms in $\bm{P}_{i}^{2}-\bm{P}_{j}^{2}\neq 0$ and inter-class distances in $(\bm{P}_{i}-\bm{P}_{j})^{2}\neq(\bm{P}_{k}-\bm{P}_{l})^{2}$ for $i\neq j$ and $\{i,j\}\neq\{k,l\}$, respectively. Note that $\lambda_{E}$ is only measured by the affine space of the class prototypes regardless of these biases. We propose to regard these factors as invariant across classes and simplify Eq. (10) as
@@
The formal proofs of Propositions 3.2, 3.3, and 3.4 are provided in Appendix A.2. The proposition reveals that when two similar samples exceed the model’s discriminability, $\mathcal{L}_{2}$ loss tends to balance the learned labels of both the harder and softer instances, instead of tending to fit a specific one of them. Notably, this objective is tailored for learning on soft-labeled data, in contrast to the calibration of one-hot-labeled samples explored in (Liang et al., 2024). In this way, we can formulate our overall objective as a simple combination of the CE and L2 losses, i.e.,
@@ 4.1 Results
We also share results on larger datasets or architectures by comparing CSM with representative methods on ImageNet with the ResNet-50 and Swin-Transformer architectures in Table 3 and Table 4, respectively. Our method performs equally or more effectively compared to these methods, especially to the mixup-based methods.
Table 3: Comparison on ImageNet using ResNet-50 architecture.
Metrics CE Mixup MbLS RegMixup RankMixup CALS CSM (Ours)
ACC 73.96 75.84 75.39 75.64 74.86 76.44 79.87
[…]
Table 4: Comparison on ImageNet using Swin-Transformer-V2 architecture.
Metrics CE LS FL FLSD MbLS CALS CSM (Ours)
ACC 75.60 75.42 75.60 74.70 77.18 77.10 81.08
[…]
We also provide qualitative comparisons using Wide-ResNet-26-10 (Zagoruyko, 2016) and DenseNet-121 (Huang et al., 2017) as the network in Appendix C, which demonstrate similar properties to those in Table 1. Meanwhile, we provide results on the Eq-Data setting in Table 2 for comparisons with non-augmentation methods for equalized training time. It is observed that our CSM can still achieve competitive or superior performance, particularly in consistently superior post-temperature ECE and AECE. Comparisons of the estimated training time is provided in Appendix C, where one can observe that the number of augmented samples per batch is the major factor affecting training time.
@@ 4.2 Discussions
1. w/o Augmented Data 77.48 12.94 12.94 2.63 2.66 1.6
2. CSM (L2) w/o Reanno. 78.92 5.20 5.20 3.33 3.24 1.2
3. CSM (L2) w/ CLIP labels. 66.60 52.78 52.78 - - -
4. CSM (L2) w/ One-hot Aug. 79.24 10.84 10.84 2.48 2.41 1.5
5. Mixup (CE) 78.47 8.68 8.68 2.14 2.19 1.3
6. Mixup (FL) 79.64 2.45 2.44 2.45 2.44 1.0
7. Mixup (L2) 78.88 2.49 2.62 2.49 2.62 1.0
8. CSM (CE) 78.93 2.47 2.14 2.47 2.14 1.0
9. CSM (FL) 77.65 2.43 2.30 2.16 2.09 0.9
10. CSM (L2, ours) 78.84 1.29 1.63 1.29 1.63 1.0
@@
Ablation Study We compare our CSM with variants in Table 6, testing different augmentation types, soft labels, and loss functions. We can observe that Mixup and CSM variants outperform the standard CE loss, improving both ECE and accuracy, highlighting the importance of data-driven network calibration. Compared with the variant w/o reannotation (or using the generation labels, Var. 2), our reannotated variant produces significantly better calibration performances, validating the accurate confidence estimation in our refined labels. Meanwhile, using the vanilla CLIP annotations (Var. 3) yields the worst ACC and pre-temperature calibration errors, primarily due to the noisy information by annotating all classes. Directly adopting class-conditioned augmentations from the diffusion model (Var. 4) can slightly rise the prediction accuracy, as also evidenced by the augmented classification literature. However, it fails to improve model calibration due to the lack of soft-labeled samples.
Among the losses used in CSM (Var. 8-10), the L2 loss is the only one that reduces ECE, AECE, and post-TS results below 2.0, suggesting it effectively balances learned confidence in augmentations. In contrast, CE and FL losses often require temperature adjustments (Var. 5,9), with CE favoring sharper labels and FL for softer ones, aligning with our theoretical expectations from Section 3.3. CSM variants, particularly with CE and L2 losses, significantly outperform traditional Mixup in terms of calibration, verifying the effectiveness of CSM’s realistic in-domain augmentations.
However, applying L2 loss to the traditional Mixup samples does not yield significantly better results (Var. 5-7). It is noteworthy that CSM uses conditioned dual-sampling of one-hot samples $\widetilde{\bm{x}}^{0}$ and $\widetilde{\bm{x}}^{1}$, which are correlated considering joint latent variable $\bm{x}^{(T)}$, while every pair of them in Mixup is sampled independently, which makes the soft-labeled samples much easier to be distinguished apart by the model. Therefore, traditional Mixup with L2 loss cannot perform noticeably better on all metrics.
Augmented Samples and Annotations To analyze the properties of augmented samples, we visualize some case images in Fig. 4. It can be seen that CSM can produce semantically sound intermediate samples compared to Mixup’s simple overlap. Meanwhile, these samples exhibit characteristics of both classes with gradual semantic change from one category to another. For images with abrupt transitions (in red numbers) or inaccurate class posteriors, our annotation paradigm yields more precise $\lambda$ values compared to the generation phase $\hat{\lambda}$.
Confidence Value Adjustment To analyze the learned confidence values, we compare the reliability diagrams of different methods across Figure 3. We examine CSM alongside three variants: 1) without augmentations, 2) with vanilla Mixup, and 3) using labels from generation. The augmentation-free variant shows severe over-confidence across the middle-to-high confidence range, while the Mixup variant also exhibits over-confidence, though less severely. Regularization-based methods that calibrate with one-hot labels can show similar over-confidence, highlighting common issues in confidence estimation. CSM with primitive labels $\hat{\lambda}$, however, is both over- and under-confident at higher and lower ranges, respectively, suggesting a misalignment between annotations and confidence levels. In contrast, CSM demonstrates precise confidence estimation across both low and high ranges, with only slight errors in the middle, confirming the effectiveness of our confidence-aware augmentation.
@@
Figure 7: Analysis of the hyperparameter $s\in[2.0,6.0]$. (a) OE and UE w.r.t. different $s$ values. (b) ECE/AECE and their post-temperature values w.r.t different $s$ values.
Table 7: Integration with Test-Time Calibration Methods.
[…]
CSM 1.29 1.63 3.16
CSM + TTA 1.39 1.53 3.15
CSM + ProCal 1.89 1.82 3.11
Integrating Test-time Methods While CSM introduces training-time augmentations for softened labels, there exists test-time augmentation methods targeting the same goal. Meanwhile, proximity problems have also been investigated in inference time. We conduct experiments to combine these methods with our CSM, as demonstrated in Table 7. Integrated with test-time augmentations (TTA), the method balances ECE and AECE effectively, achieving an optimized AECE of 1.53 on CIFAR-100. Compared to (Hekler et al., 2023) using test-time sample-wise scaling, CSM employs training-time augmentation with inter-sample augmentations to expand the proximity space, enhancing calibration robustness. Combined with proximity method ProCal, we find that the proximity-informed metric PIECE displays better results, which validates the robustness growth related to proximity from the integration. Despite these growths, the overall calibration improvement is relatively small, indicating the effective sample augmentation and debiased proximity learning of our method.